Someone please help
Find the midway point of the line segment from A(-2,9) to B(4,5)
possible answers:
c(1,7)
D(3,7)
P(4,9)
Q(5,9)
please show me how to do this also.
Thank you

Answer by bucky(2189)

Find the midway point of the line segment from A(-2,9) to B(4,5)
The coordinates are found by finding the halfway point between the x values of the points
and then finding the halfway point between the y values.
.
First the x values. One x value is at -2 and the other at +4. There are 6 units between
these numbers. So half the distance between them is 3 units. Either add 3 units to -2
or subtract 3 units from 4 to get to the point +1 on the number line. If you prefer
you can just average the x points. Add -2 and +4 and divide by 2. You will get the same
result ... -2 + 4 = +2 and divide that by 2 gives you +1 as the x coordinate of the
midway point. Note that from your answer list you have only one answer with the x value
of +1, so you already know the answer. But let's continue with finding the y value
just for practice.
.
The y values of the two points are 9 and 5. The length of the y distance between these
points is 4. We can split the difference ... 4/2 = 2 and then either add this 2 to 5 or
subtract the 2 from 9 and get the answer as y = +7. Or we could average the 9 and 5
by adding them together and dividing by 2 ... (9 + 5)/2 = 14/2 = +7.
.
The coordinates that we have found are (1,7) and that definitely is in your list of answers.
.
Not too tough once you think about this procedure and practice it a little.
.
Hope this helps make math a little easier for you.